# V2V-LLM: Vehicle-to-Vehicle Cooperative Autonomous Driving with Multi-Modal Large Language Models

### v1

@@ 3.4 Evaluation Metrics
Following prior works [39, 43], we evaluate the performance of different models on the grounding questions (Q1, Q2, Q3) and the notable object identification question (Q4) with the F1 score, precision, and recall. For the planning question (Q5), the evaluation metrics are L2 errors and collision rates.
@@ 4 V2V-LLM
In addition to the proposed dataset, we propose a competitive baseline model, V2V-LLM, for this LLM-based collaborative driving problem, as shown in Fig. 3. This model is a Multi-modal LLM (MLLM) that takes the individual perception features of every CAV as the vision input, a question as the language input, and generates an answer as the language output.
@@ 4.1 LiDAR-based Input Features
4.2 LiDAR-based LLM
@@
Training: Our V2V-LLM uses LLaVA-v1.5-7b [25]’s Vicuna [7] as the LLM backbone. To train our model, we initialize it by loading the pre-trained LLaVA-v1.5-7b [25]’s checkpoint. We freeze the LLM and the point cloud feature encoder, and only finetune the projector and the LoRA [14] parts of the model. During training, we use batch size 32. We train our Q1 for 1 epoch, and other QA types for 10 epochs. For all other training settings and hyperparameters, we use the same ones from LLaVA-v1.5-7b [25].
@@ 5.1 Baseline Methods
We follow V2V4Real [53] and V2X-Real [48] to establish a benchmark for our proposed V2V-QA dataset with experiments on baseline methods using different fusion approaches: no fusion, early fusion, intermediate fusion, and our proposed baseline, LLM fusion (Fig. 3). Compared with our method, other baseline methods also adopt the same projector and LLM architecture but with different point cloud feature encoders. Their feature encoder diagrams can be seen in Fig. 4.
@@
No fusion: Only a single CAV’s LiDAR point cloud is fed to a single 3D object detector to extract the scene-level feature map and the object-level feature vectors, which are then used as the LLM’s visual input. The performance is expected to be worse than all other cooperative perception approaches because the other CAV’s sensor input is ignored.
@@
Early fusion: The LiDAR point cloud from two CAVs is merged first. Then the merged point cloud is used as input to a 3D object detector to extract the visual features as the visual input to the LLM. This approach may achieve good performance by using all the raw sensor input but requires much higher communication bandwidth when compared with other fusion approaches. This limitation causes early-fusion approaches to be impractical for deployment on real-world autonomous vehicles.
@@
Intermediate fusion: Prior research CoBEVT [50], V2X-ViT [51], and AttFuse [52] propose different cooperative detection models that can merge feature maps from multiple CAVs via attention mechanisms. Such approaches require less communication bandwidth and can still achieve good performance. In our benchmark, we extract the features from those cooperative detection models as the visual input to the LLM.
@@
LLM fusion: We categorize our proposed V2V-LLM as a new type of fusion method, LLM fusion, which lets each CAV perform its individual 3D object detection to extract the scene-level feature maps and object-level feature vectors, and uses the LLM to fuse the features from multiple CAVs. This approach is related to the traditional late fusion method that performs individual 3D object detection and aggregates the results by non-maximum suppression (NMS). Instead of applying NMS, our method adopts LLM to perform more tasks than just detection.
@@
V2V-LLM (Ours) 70.0 80.1 62.2 30.8 36.3 26.7 21.2 21.5 20.8 40.7 59.7 61.9 57.6 4.99 3.00 0.203
Table 3: V2V-LLM’s performance in V2V-QA’s testing split in comparison with baseline methods. Q1: Grounding at a reference location. Q2: Grounding behind a reference object at a location. Q3: Grounding behind a reference object in a direction. QGr: Average of grounding (Q1, Q2, and Q3). Q4: Notable object identification. Q5: Planning. P: Precision. R: Recall. L2: L2 distance error. CR: Collision Rate. Comm: Communication cost. In each column, the best results are in boldface, and the second-best results are in underline.
Method L2 (m) CR (%) Comm (MB) $\downarrow$
No Fusion 3.84 6.52 9.30 6.55 1.31 4.76 7.63 4.57 0
Early Fusion 3.68 6.19 8.74 6.20 0.96 3.86 5.83 3.55 0.96
Intermediate Fusion
AttFuse [52] 4.06 6.78 9.64 6.83 1.42 4.41 6.53 4.12 0.20
V2X-ViT [51] 4.21 7.05 9.99 7.08 1.33 4.82 6.85 4.33 0.20
CoBEVT [50] 3.97 6.71 9.47 6.72 0.93 3.74 6.96 3.88 0.20
LLM Fusion
V2V-LLM (ours) 2.96 4.97 7.05 4.99 0.55 3.19 5.25 3.00 0.203
Table 4: V2V-LLM’s planning performance in V2V-QA’s testing split in comparison with baseline methods. L2: L2 distance error. CR: Collision Rate. In each column, the best results are in boldface. and the second-best results are in underline.
5.2 Experimental Results
5.2.1 Grounding
Our V2V-LLM and baseline methods’ performance on V2V-QA’s $3$ types of grounding questions can be seen in Tab. 3. We can see that all fusion methods outperform the no-fusion approach. Early fusion, V2X-ViT [50] intermediate fusion, and our proposed V2V-LLM achieve best results in Q1, Q2, and Q3 respectively. In average, CoBEVT [50] achieves the best result and V2V-LLM  achieves the second-best result. Such results indicate that our MLLM also has a promising capability of fusing scene-level feature maps and object-level feature vectors from multiple CAVs. And its performance is not too far away from the performance of other, specially designed intermediate fusion models.
5.2.2 Notable Object Identification
Tab. 3 shows the performance on the notable object identification task (Q4). Our proposed V2V-LLM outperforms other methods. Compared with the aforementioned grounding tasks, this notable object identification task requires more spatial understanding and reasoning ability to identify the objects close to the planned future waypoints provided by the asking CAV. For such a task, our V2V-LLM, which lets the LLM perform both feature fusion and question answering, achieves the best results.
5.2.3 Planning
Tab. 4 shows the detailed performance on the planning task (Q5). Our proposed V2V-LLM outperforms other methods in this most important question which requires more spatial understanding and reasoning ability to generate a safe future trajectory that can avoid potential collisions.
5.2.4 Summary
Overall, V2V-LLM achieves the best results in the notable object identification and planning tasks, which are more important than the grounding tasks in autonomous driving applications. V2V-LLM also achieves the second-best result in the grounding tasks. In terms of communication costs, V2V-LLM shares both scene-level-feature maps and object-level-feature vectors and only increases communication costs by $1.5\%$ in comparison to other intermediate fusion baseline methods.
@@ 5.2.4 Summary
Scene-level only 69.9 74.9 65.5 15.4 19.9 12.6 17.9 26.9 13.5 34.4 43.2 40.2 46.7 7.21 15.55 0.20
Object-level only 69.0 80.9 60.1 26.9 34.7 21.9 17.6 18.3 16.9 37.8 52.6 57.3 48.6 5.24 7.78 0.003
V2V-LLM (ours) 70.0 80.1 62.2 30.8 36.3 26.7 21.2 21.5 20.8 40.7 59.7 61.9 57.6 4.99 3.00 0.203
@@
Table 5: Ablation study in V2V-QA’s testing split. Q1: Grounding at a reference location. Q2: Grounding behind a reference object at a location. Q3: Grounding behind a reference object in a direction. QGr: Average of grounding (Q1, Q2, and Q3). Q4: Notable object identification. Q5: Planning. P: Precision. R: Recall. L2: L2 distance error. CR: Collision Rate. Comm: Communication cost.
5.3 Ablation Study
We experiment with variants of our V2V-LLM model that use either only the scene-level feature maps or only the object-level feature vectors as the visual input. The ablation results can be seen in Tab. 5. Overall, we can observe that both scene-level feature maps and object-level feature vectors contribute to final performance in all QA tasks. In Q1 and Q3, the scene-level-only model and object-level-only model achieve similar performance. For other QA tasks, the object-level-only model outperforms the scene-level-only model by a large margin. This implies that the object-level features are easier for LLM to digest, which is consistent with the results observed in the previous work with the TOKEN model [39].
Figure 5: V2V-LLM’s grounding results on V2V-QA’s testing split. Magenta $\circ$: reference locations in questions. Yellow $+$: model output locations. Green $\circ$: ground-truth answers.
Figure 6: V2V-LLM’s notable object identification and planning results on V2V-QA’s testing split. For notable object identification, Magenta curve: planned future trajectories in questions. Green $\circ$: ground-truth notable object locations. Yellow $+$ and Cyan $\times$: model identification outputs corresponding to CAV_EGO and CAV_1, respectively. For planning, Green line: future trajectories in ground-truth answers. Yellow curve and Cyan curve: model planning outputs corresponding to CAV_EGO and CAV_1, respectively.
5.4 Qualitative Results
Fig. 5 shows our V2V-LLM’s grounding results and the ground truth with visualization on V2V-QA’s testing split. We can observe that our V2V-LLM  is able to locate the objects given the provided reference information for each of the $3$ types of grounding questions. Fig. 6’s left part shows our V2V-LLM’s notable object identification results. V2V-LLM  demonstrate its capability of identifying multiple objects near the planned future trajectories specified in the questions for each CAV. Fig. 6’s right part shows V2V-LLM’s planning results. Our model is able to suggest future trajectories that avoid potential collisions with nearby objects. Overall, the outputs of our model closely align with the ground-truth answers across all question types, indicating its robustness in cooperative autonomous driving tasks.
6 Conclusion
In this work, we expand the research scope of cooperative autonomous driving by integrating the use of LLM-based methods, aimed at improving the safety of future autonomous driving systems. We propose a new problem setting and create a novel V2V-QA dataset and benchmark that includes grounding, notable object identification, and planning question-answering tasks designed for varieties of cooperative driving scenarios. We propose a baseline model V2V-LLM that fuses each CAV’s individual perception information and performs visual and language understanding to answer driving-related questions from any CAV. In comparison to other baseline methods adopted from state-of-the-art cooperative perception algorithms, our proposed V2V-LLM achieves comparative performance in the grounding tasks and outperforms all other baseline methods in the more important notable object identification and planning tasks. These experimental results indicate that V2V-LLM is promising as a unified foundation model that can effectively perform perception and planning tasks for cooperative autonomous driving. We believe our V2V-QA  dataset will bring the cooperative driving research field to the next stage.
7 Acknowledgement
Supplementary Material
8 Additional Planning Results

### v3

@@ 3.4 Evaluation Metrics
We follow prior works [39, 43]’s approach to evaluate model performance. For the grounding questions (Q1, Q2, Q3) and the notable object identification question (Q4), the evaluation metrics are F1 score, precision, and recall. The ground-truth answers and model outputs contain objects’ center locations. If the center distance between the ground-truth answer and the model output is less than a threshold value, this output is considered as a true positive. We set the threshold value to be $4$ meters, a typical length of a vehicle.
For the planning question (Q5), the evaluation metrics are L2 errors and collision rates. The ground-truth answers and model outputs contain $6$ future waypoints, so we calculate the L2 errors on those waypoints. When calculating collision rates, we assume each of the CAVs’ bounding box sizes is $4$ meters in length, $2$ meters in width, and $1.5$ meters in height. We place each CAV’s bounding box at the model’s output future waypoints and calculate the Intersection-over-Union (IOU) between the CAV’s bounding box and every ground-truth object’s annotation bounding box in those future frames. If the IOU is larger than 00, it is considered as a collision.
@@ 4 V2V-LLM
We also propose a competitive baseline model, V2V-LLM, for this LLM-based collaborative driving problem, as shown in Fig. 3. Our model is a Multi-Modal LLM (MLLM) that takes the individual perception features of every CAV as the vision input, a question as the language input, and generates an answer as the language output.
@@ 4.1 LiDAR-based Input Features
4.2 LiDAR-based Multi-Modal LLM
@@
Training: We use 8 NVIDIA A100-80GB GPUs to train our model. Our V2V-LLM uses LLaVA-v1.5-7b [25]’s Vicuna [7] as the LLM backbone. To train our model, we initialize it by loading the pre-trained LLaVA-v1.5-7b [25]’s checkpoint. We freeze the LLM and the point cloud feature encoder, and finetune the projector and the LoRA [14] parts of the model. During training, we use batch size 32. Adam optimizer is adopted for training with a starting learning rate $52e-$ 5 and a cosine learning rate scheduler with a 3% warm-up ratio. For all other training settings and hyperparameters, we use the same ones from LLaVA-v1.5-7b [25].
@@ 5.1 Baseline Methods
We follow V2V4Real [53] and V2X-Real [48] to establish a benchmark for our proposed V2V-QA dataset with experiments on baseline methods using different fusion approaches: no fusion, early fusion, intermediate fusion, and our proposed baseline, LLM fusion (Fig. 3). The baseline methods also adopt the same projector and LLM architecture as in our V2V-LLM but with different point cloud feature encoders, as shown in Fig. 4. In some driving sequences of V2X-split that have point clouds from roadside infrastructures, we also include them as perception input in the same way as using CAVs’ point clouds.
@@
No fusion: Only a single CAV’s LiDAR point cloud is fed to a single 3D object detector to extract the scene-level and object-level features as the LLM’s visual input. The performance is expected to be worse than all other cooperative perception approaches.
@@
Early fusion: The LiDAR point cloud from two CAVs is merged first. Then the merged point cloud is used as input to a 3D object detector to extract the visual features as the visual input to the LLM. This approach requires much higher communication cost and is less practical for deployment on real-world autonomous vehicles.
@@
Intermediate fusion: Prior work CoBEVT [50], V2X-ViT [51], and AttFuse [52] propose cooperative detection models that merge feature maps from multiple CAVs via attention mechanisms. Such approaches require less communication cost and can still achieve good performance. In our benchmark, we extract the features from those cooperative detection models as the input tokens to the LLM.
@@
LLM fusion: We categorize our proposed V2V-LLM as a new type of fusion method, LLM fusion, which lets each CAV perform its individual 3D object detection to extract the scene-level feature maps and object-level feature vectors, and uses the Multi-Modal LLM to fuse the features from multiple CAVs. This approach is related to the traditional late fusion method that performs individual 3D object detection and aggregates the results by non-maximum suppression (NMS). Instead of applying NMS, our method adopts LLM to perform more tasks than just detection.
Method V2V-split V2X-split Comm(MB) $\downarrow$
Q1 Q2 Q3 QGr Q4 Q5 Q1 Q2 Q3 QGr Q4 Q5
F1 $\uparrow$ F1 $\uparrow$ F1 $\uparrow$ F1 $\uparrow$ F1 $\uparrow$ L2 (m) $\downarrow$ CR (%) $\downarrow$ F1 $\uparrow$ F1 $\uparrow$ F1 $\uparrow$ F1 $\uparrow$ F1 $\uparrow$ L2 (m) $\downarrow$ CR (%) $\downarrow$
No Fusion 66.6 22.6 17.2 35.5 47.3 6.55 4.57 55.7 21.4 25.2 34.1 64.4 2.31 9.21 0
Early Fusion 73.5 23.3 20.8 39.2 53.9 6.20 3.55 59.7 23.3 26.1 36.4 67.6 2.12 8.61 1.9208
Intermediate Fusion
AttFuse [52] 70.7 26.4 18.4 38.5 56.9 6.83 4.12 58.9 23.9 26.3 36.4 65.9 2.19 8.39 0.4008
V2X-ViT [51] 70.8 28.0 22.6 40.5 57.6 7.08 4.33 59.6 24.2 26.1 36.6 65.0 2.29 8.86 0.4008
CoBEVT [50] 72.2 29.3 21.3 40.9 57.6 6.72 3.88 - - - - - - - 0.4008
LLM Fusion
V2V-LLM (Ours) 70.0 30.8 21.2 40.7 59.7 4.99 3.00 60.5 25.3 26.7 37.5 69.3 1.71 6.89 0.4068
Table 3: V2V-LLM’s testing performance in V2V-QA’s V2V-split and V2X-split in comparison with baseline methods. Q1: Grounding at a reference location. Q2: Grounding behind a reference object at a location. Q3: Grounding behind a reference object in a direction. QGr: Average of grounding (Q1, Q2, and Q3). Q4: Notable object identification. Q5: Planning. L2: L2 distance error. CR: Collision rate. Comm: Communication cost. In each column, the best results are in boldface, and the second-best results are in underline. More detailed performance evaluation can be seen in the supplementary material.
5.2 Experimental Results
5.2.1 Grounding
Our V2V-LLM and baseline methods’ performance on V2V-QA’s $3$ types of grounding questions can be seen in Tab. 3 for V2V-split and V2X-split, respectively. CoBEVT [50] is not included in V2X-split’s result because V2X-Real [48] does not release its CoBEVT [50] baseline model. In average, V2V-LLM  achieves similar performance in V2V-split and outperforms all other baseline methods in V2X-split. Such results indicate that our V2V-LLM has a promising capability of fusing perception features from multiple CAVs to answer grounding questions.
5.2.2 Notable Object Identification
Tab. 3 show the performance on the notable object identification task (Q4). Our proposed V2V-LLM outperforms all other methods in both V2V-split and V2X-split. Compared with the aforementioned grounding tasks, this notable object identification task requires more spatial understanding ability to identify the objects close to the planned future waypoints. For such a task, our V2V-LLM, which lets the Multi-Modal LLM perform both perception feature fusion and question answering, achieves the best results.
5.2.3 Planning
Tab. 3 show the performance of the planning task (Q5) for V2V-split and V2X-split, respectively. Our proposed V2V-LLM outperforms other methods in this safety-critical task to generate a future trajectory that aims to avoid potential collisions. More planning performance evaluation can be seen in the supplementary material.
5.2.4 Communication Cost and Scaling Analysis
In our centralized setting, each CAV sends one scene-level feature map ($\leq 0.2$MB), one set of individual object detection result parameters ($\leq 0.003$MB), one question ($\leq 0.0002$MB) to the LLM computing node and receives one answer ($\leq 0.0002$MB) at each timestep. If there are $N_{v}$ CAVs and each asks $N_{q}$ questions, the communication cost of each CAV is $(0.2+0.003+(0.0002+0.0002)N_{q})=(0.203+0.0004N_{q})$ MB, and the cost of the LLM is $(0.2+0.003+(0.0002+0.0002)N_{q})N_{v}=(0.203N_{v}+0.0004N_{q}N_{v})$ MB, as shown in Tab. 4. Note that each CAV only needs to send the same features to the LLM computing node once at each timestep because the LLM node can save and reuse them to answer multiple questions from the same or different CAVs at the same timestep. Detailed scaling analysis can be seen in the supplementary material.
Setting Each CAV Centralized LLM
Centralized $0.203+0.0004N_{q}$ $0.203N_{v}+0.0004N_{q}N_{v}$
Table 4: Communication cost (MB) and scaling analysis. $N_{v}$: number of CAVs. $N_{q}$: number of questions asked by each CAV at each timestep.
5.2.5 Summary
Overall, V2V-LLM achieves the best results in the notable object identification and planning tasks, which are critical in autonomous driving applications. V2V-LLM also achieves competitive results in the grounding tasks. In terms of communication costs, V2V-LLM only increases communication costs by $1.5\%$ in comparison to other intermediate fusion baseline methods. More detailed evaluation and analysis can be seen in the supplementary material.
Method Q1 Q2 Q3 QGr Q4 Q5
F1 $\uparrow$ F1 $\uparrow$ F1 $\uparrow$ F1 $\uparrow$ F1 $\uparrow$ L2 (m) $\downarrow$ CR (%) $\downarrow$
Scene only 69.9 15.4 17.9 34.4 43.2 7.21 15.55
Object only 69.0 26.9 17.6 37.8 52.6 5.24 7.78
Scratch 67.6 26.5 17.2 37.1 49.3 6.30 5.01
V2V-LLM 70.0 30.8 21.2 40.7 59.7 4.99 3.00
Table 5: Ablation study in V2V-QA’s V2V-split testing set.
5.3 Ablation Study
Input Features: We experiment with variants of our V2V-LLM model that use either only the scene-level feature maps or only the object-level feature vectors as the visual input. The ablation results can be seen in Tab. 5. Both types of features contribute to final performance in all QA tasks. In general, the object-level-only model outperforms the scene-level-only model. This implies that the object-level features are easier for LLM to digest, which is consistent with the results observed in the previous work with the TOKEN model [39].
Training from Scratch: Tab. 5 shows that training from scratch achieves worse performance, meaning that pre-training with LLaVA’s VQA tasks improves our V2V-LLM’s performance in V2V-QA. More detailed ablation results can be seen in the supplementary material.
Figure 5: V2V-LLM’s grounding results on V2V-QA’s testing set. Magenta $\circ$: reference locations in questions. Yellow $+$: model output locations. Green $\circ$: ground-truth answers.
Figure 6: V2V-LLM’s notable object identification and planning results on V2V-QA’s testing set. For notable object identification, Magenta curve: planned future trajectories in questions. Green $\circ$: ground-truth notable object locations. Yellow $+$ and Cyan $\times$: model identification outputs corresponding to CAV_EGO and CAV_1, respectively. For planning, Green line: future trajectories in ground-truth answers. Yellow curve and Cyan curve: model planning outputs corresponding to CAV_EGO and CAV_1, respectively.
5.4 Qualitative Results
Fig. 5 shows our V2V-LLM’s grounding results and the ground truth with visualization on V2V-QA’s testing set. We can observe that our V2V-LLM  is able to locate the objects given the provided reference information for each of the $3$ types of grounding questions. Fig. 6’s left part shows our V2V-LLM’s notable object identification results. V2V-LLM  demonstrate its capability of identifying multiple objects near the planned future trajectories specified in the questions for each CAV. Fig. 6’s right part shows V2V-LLM’s planning results. Our model is able to suggest future trajectories that avoid potential collisions with nearby objects. Overall, the outputs of our model closely align with the ground-truth answers across all question types, indicating its robustness in cooperative autonomous driving tasks.
6 Conclusion
In this work, we expand the research scope of cooperative autonomous driving by integrating the use of Multi-Modal LLM-based methods, aimed at improving the safety of future autonomous driving systems. We propose a new problem setting and create a novel V2V-QA dataset and benchmark that includes grounding, notable object identification, and planning question-answering tasks designed for varieties of cooperative driving scenarios. We propose a baseline model V2V-LLM that fuses each CAV’s individual perception information and performs visual and language understanding to answer driving-related questions from any CAV. Our proposed V2V-LLM outperforms all other baselines adopted from state-of-the-art cooperative perception algorithms in the notable object identification and planning tasks, and achieves competitive performance in the grounding tasks. These experimental results indicate that V2V-LLM is promising as a unified multi-modal foundation model that can effectively perform perception and planning tasks for cooperative autonomous driving. We will publicly release our V2V-QA  dataset and code to facilitate the open-source research, and believe it will bring the cooperative driving research to the next stage.
7 Acknowledgement
Supplementary Material
8 Detailed Evaluation Results
Tabs. 6 and 7 summarize the detailed evaluation results of our V2V-LLM and other baseline methods in V2V-QA’s V2V-split and V2X-split. In addition, Tabs. 8 and 9 show the detailed planning performance. For the grounding task, our V2V-LLM achieves competitive results in V2V-split and outperforms all other baseline methods in V2X-split. More importantly, for the notable object identification task and the planning task, our V2V-LLM outperforms all other baseline methods in both V2V-split and V2X-split.
@@ 8 Detailed Evaluation Results
V2V-LLM (Ours) 70.0 80.1 62.2 30.8 36.3 26.7 21.2 21.5 20.8 40.7 59.7 61.9 57.6 4.99 3.00 0.4068
Table 6: V2V-LLM’s testing performance in V2V-QA’s V2V-split in comparison with baseline methods. Q1: Grounding at a reference location. Q2: Grounding behind a reference object at a location. Q3: Grounding behind a reference object in a direction. QGr: Average of grounding (Q1, Q2, and Q3). Q4: Notable object identification. Q5: Planning. P: Precision. R: Recall. L2: L2 distance error. CR: Collision rate. Comm: Communication cost. In each column, the best results are in boldface, and the second-best results are in underline.
@@
No Fusion 55.7 71.6 45.5 21.4 33.2 15.8 25.2 26.2 24.2 34.1 64.4 66.1 62.7 2.31 9.21 0
Early Fusion 59.7 70.6 51.8 23.3 34.0 17.7 26.1 28.0 24.5 36.4 67.6 69.3 66.0 2.12 8.61 1.9208
Intermediate Fusion
AttFuse [52] 58.9 71.1 50.3 23.9 34.3 18.4 26.3 28.3 24.6 36.4 65.9 67.0 64.9 2.19 8.39 0.4008
V2X-ViT [51] 59.6 69.6 52.1 24.2 33.2 19.1 26.1 28.2 24.3 36.6 65.0 64.8 65.3 2.29 8.86 0.4008
LLM Fusion
V2V-LLM (Ours) 60.5 69.5 53.6 25.3 34.9 19.8 26.7 27.0 26.4 37.5 69.3 71.9 66.8 1.71 6.89 0.4068
@@
Table 7: V2V-LLM’s testing performance in V2V-QA’s V2X-split in comparison with baseline methods. Q1: Grounding at a reference location. Q2: Grounding behind a reference object at a location. Q3: Grounding behind a reference object in a direction. QGr: Average of grounding (Q1, Q2, and Q3). Q4: Notable object identification. Q5: Planning. P: Precision. R: Recall. L2: L2 distance error. CR: Collision rate. Comm: Communication cost. In each column, the best results are in boldface, and the second-best results are in underline.
Method L2 (m) CR (%)
No Fusion 3.84 6.52 9.30 6.55 1.31 4.76 7.63 4.57
Early Fusion 3.68 6.19 8.74 6.20 0.96 3.86 5.83 3.55
Intermediate Fusion
AttFuse [52] 4.06 6.78 9.64 6.83 1.42 4.41 6.53 4.12
V2X-ViT [51] 4.21 7.05 9.99 7.08 1.33 4.82 6.85 4.33
CoBEVT [50] 3.97 6.71 9.47 6.72 0.93 3.74 6.96 3.88
LLM Fusion
V2V-LLM (ours) 2.96 4.97 7.05 4.99 0.55 3.19 5.25 3.00
Table 8: V2V-LLM’s planning performance in V2V-QA’s V2V-split in comparison with baseline methods. L2: L2 distance error. CR: Collision rate. In each column, the best results are in boldface, and the second-best results are in underline.
Method L2 (m) CR (%)
No Fusion 1.33 2.28 3.31 2.31 2.52 9.54 15.57 9.21
Early Fusion 1.24 2.10 3.00 2.12 3.51 8.37 13.93 8.61
Intermediate Fusion
AttFuse [52] 1.27 2.17 3.11 2.19 2.40 9.07 13.70 8.39
V2X-ViT [51] 1.34 2.27 3.25 2.29 1.41 9.89 15.28 8.86
LLM Fusion
V2V-LLM (ours) 0.99 1.70 2.45 1.71 2.17 6.79 11.71 6.89
Table 9: V2V-LLM’s planning performance in V2V-QA’s V2X-split in comparison with baseline methods. L2: L2 distance error. CR: Collision rate. In each column, the best results are in boldface, and the second-best results are in underline.
9 Detailed Communication Cost and Scaling Analysis
In our centralized setting, assume that there is one centralized LLM computing node, $N_{v}$ CAVs, and each CAV asks $N_{q}$ questions at each timestep. Each CAV sends one scene-level feature map ($\leq 0.2$MB), one set of individual object detection result parameters ($\leq 0.003$MB), $N_{q}$ questions (each $\leq 0.0002$MB) to the LLM and receives $N_{q}$ answers (each $\leq 0.0002$MB) at each timestep. Note that each CAV only needs to send the same features to the LLM once at each timestep because the LLM node can save and reuse them to answer multiple questions from the same or different CAVs at the same timestep. The communication cost of each CAV is: $0.2+0.003+(0.0002+0.0002)N_{q}=(0.203+0.0004N_{q})$ MB. The LLM receives $N_{v}$ scene-level feature maps, $N_{v}$ set of individual object detection result parameters, $N_{q}N_{v}$ questions and returns $N_{q}N_{v}$ answers. The communication cost of the centralized LLM is $(0.2+0.003+(0.0002+0.0002)N_{q})N_{v}=(0.203N_{v}+0.0004N_{q}N_{v})$ MB.
Alternatively, one can also consider a decentralized setting that deploys one LLM in each CAV. In this setting, each CAV receives the features from all other CAVs and does not need to send or receive any questions or answers. The communication cost of each CAV is $(0.2+0.003)(N_{v}-1)=0.203(N_{v}-1)$ MB. Tab. 10 summarizes the communication cost and scaling analysis in the aforementioned settings. There could be more different decentralized settings. Which setting works best in terms of communication costs is beyond the current focus of our work.
Setting Each CAV Centralized LLM
Centralized $0.203+0.0004N_{q}$ $0.203N_{v}+0.0004N_{q}N_{v}$
Decentralized $0.203(N_{v}-1)$
Table 10: Communication cost (MB) and scaling analysis. $N_{v}$: number of CAVs. $N_{q}$: number of questions asked by each CAV at each timestep.
10 Planning Results with Temporal Inputs